# Automate Strategy Finding with LLM in Quant investment

### v1

@@ 1. Introduction
According to the survey by Grand View Research, in 2023 the global alternative data market was valued at 7.2bn USD and is expected to expand at the compound annual growth rate (CAGR) of 52.1% from 2023 to 2030 representing a complex and dynamic ecosystem that attracts millions of investors seeking profitable investment opportunities. (vladi2024quantitative, ) Recent advancements in quantitative trading have significantly enhanced the ability to analyze and process vast amounts of financial data, leading to the development of sophisticated trading strategies. One of the key areas of focus in quantitative trading is Alpha mining—the process of discovering and refining predictive signals, or alpha factors, that forecast asset returns. (tulchinsky2015finding, ) Eugene Fama proposed the Efficient Market Hypothesis (EMH) to present that stock price will be the reflection of all the market available information. (fama1970efficient, ) However with the market condition change the same information may make different influence on the market, some researchers also found information asymmetry. (allen1999genetic, ) Here we found three major challenges to the current Alpha mining process (1) Rigidness of traditional methods: traditional methods for identifying alpha factors in finance often rely on heuristic rules and financial expertise. These rule-based approaches, typically perform well only under specific market conditions and lack the flexibility to adapt to varying market dynamics. (2) Data diversity and integration, recent research has explored the application of machine learning techniques to uncover alpha factors from diverse data sources, such as economic news, social media trends, and historical price movements. (cui2021alphaevolve, ; allen1999genetic, ; yu2023generating, ) (3) Adapting to market variability, Market dynamics are fluid, and strategies that perform well in one environment may fail in another. In recent years many researchers proposed to use deep learning methods to predict the market (xu2018stock, ) and form their strategy. (chen2019investment, ) Efficiently mining and utilizing alpha factors across different market conditions remains a significant challenge.
@@
To address this challenge, we propose a novel framework that leverages the exploratory power of Large Language Models (LLMs) for Alpha mining and strategy optimization. (1) Flexible Alpha mining with LLMs, the framework starts with the collection and categorization of potential alpha factors from academic research, financial literature, and proprietary datasets. These seed alphas are filtered and categorized using the potential of the LLM, which ensures a comprehensive and diverse set of candidate seed Alphas for further research. All these seed alpha factors are well categorized as momentum factors, fundamental factors, liquidity factors, etc., and based on traditional finance research we believe that these factors are independent of each other. (xu2018stock, ) Once new research work is published, our model will be able to incremental update the seed alphas factory. This dynamic approach allows for the discovery of adaptable alpha factors, overcoming the rigidity of traditional methods. (2) Multi-Agent multimodal market evaluation, the seed Alpha factors undergo rigorous back-testing against historical market data to assess their performance. This includes back test under different market conditions to determine which factors can respond to the market for excess returns under different market conditions. To handle the variability in market conditions, we employ a multi-Agent approach. This method allows specialized agents to evaluate the relevance and effectiveness of each alpha factor under multimodal market scenarios. The multi-Agent architecture captures the nuanced dynamics of various market environments from different perspectives. Based on the current market condition the multi-Agent part can filter the out performance seed alphas for us further strategy design. (3) Dynamic strategy optimization, following the evaluation and back-testing phases, a weight gating layer assigns optimal weights to each seed alpha based on current market conditions. This dynamic weighting mechanism ensures that the strategy adapts to changing market environments, optimizing overall performance. The weighted alpha strategies are then aggregated to form a comprehensive investment strategy that maximizes returns while managing risk. Our approach integrates state-of-the-art techniques in machine learning and finance to develop a robust system capable of identifying and optimizing alpha strategies across multiple asset classes. As Figure 7 shows here the work inspired from the idea in real-life quant company practice, for the real quant investment company they generally have four major departments. The first one is the research department’s response to find the valuable trading signals and record them; the second part is the trading department need to analyze the current market condition and select the ideal alphas for future portfolio construction; the third part is the risk management part, they need to evaluate and make the strategy suitable for the risk preference; the last part will be the portfolio manager, they mix up the traders idea construct the final trading strategy for profit.(fabozzi2010quantitative, )
@@ 2.2. Seed Alpha Mining & Selection
In quantitative trading, the static nature of traditional heuristic and rule-based alpha mining methods poses significant limitations (e.g. Information coefficient), particularly in the context of rapidly changing market conditions. (zhang2020information, ) These methods often fail to adapt dynamically, leading to suboptimal performance and the selected Alpha may loss function. Moreover, the reliance on structured financial data, such as historical prices and fundamental indicators, overlooks the rich insights that can be derived from alternative data sources like economic news and social media trends. To address these challenges, we propose a novel framework that leverages the exploratory power of Large Language Models (LLMs) and a multi-agent system for seed alpha mining and selection. Our approach integrates state-of-the-art machine learning techniques with financial principles to create a robust system capable of dynamically adjusting to market variability and optimizing alpha strategies across $k$ categories of seed alphas. This multimodal framework links current market conditions with alpha back-test results, allowing LLMs to identify the most suitable seed alphas while assigning confidence scores $\theta$ to ensure reliability. Specialized agents further customize the evaluation based on risk preferences and expertise, ensuring a comprehensive and adaptive solution for maximizing returns and managing risks in diverse market environments.
@@ 3.3. Multimodal & multi-Agent Evaluate Alpha set
The multi-agent system comprises various agents, each designed with distinct risk preferences and investment strategies. These agents analyze the multimodal data and evaluate the seed alphas based on their respective criteria. Each agent conducts a risk preference analysis, assessing the seed alphas according to its predefined risk parameters. This ensures that the evaluation covers a wide spectrum of risk perspectives. Agents then select alpha factors that align with their risk preferences and market analyses, ensuring relevance and robustness under different market scenarios. To enhance the reliability of the selected alpha factors, a confidence scoring mechanism is implemented. Each agent assigns a confidence score to the alpha factors based on their analysis and historical performance, reflecting the agent’s assessment of the alpha factor’s predictive power and robustness. Selected alpha factors undergo rigorous backtesting using historical market data to evaluate their performance under various market conditions. This step is crucial for identifying the most promising alpha factors for the final strategy. Key evaluation metrics include performance indicators such as the Information Coefficient (IC) and Sharpe Ratio etc.(sharpe1994sharpe, ) Additionally, seed alpha factors from each category must meet specific confidence score thresholds to be considered for the final selection, ensuring that only the most reliable and robust alpha factors are chosen.
@@ 3.4. Optimal weight alpha strategy
To optimize the weights of selected seed alphas, we employ a Deep Neural Network (DNN) to fit alpha values to future yields. The network architecture comprises three layers: an input layer, a hidden layer, and an output layer. The input layer takes daily alpha calculations derived from historical closing prices as input features. The hidden layer, equipped with ten nodes, utilizes ReLU activation functions to introduce non-linearity and enhance the model’s learning capability. (chen2020dynamic, ) The output layer, consisting of a single node, predicts future yields. During the training phase, the network employs backpropagation and gradient descent to minimize the loss function, which quantifies the discrepancy between predicted and actual yields. This iterative process adjusts the network’s weights and biases to improve prediction accuracy. To ensure the model’s generalizability and prevent overfitting, a separate validation set is utilized. This methodology establishes a robust framework for predicting future yields based on historical alpha values, thus forming an effective investment strategy. The DNN model processes the input data through the hidden layer, transforming it with learned weights and biases. The final output is generated by applying another set of weights, biases, and an activation function in the output layer. Our approach demonstrates the efficacy of leveraging deep learning techniques to optimize alpha weights, thereby enhancing the predictive performance of investment strategies.
@@ 4.1. Datasets
Our study focuses on financial data from the Chinese A-shares market, specifically targeting the SSE 50 Index. The dataset encompasses six primary features as original inputs for our Alpha factors: open, close, high, low, volume, and volume-weighted average price (VWAP). To ensure rigorous evaluation and robust model performance, the dataset is divided into distinct time periods: Training Set: January 1, 2021, to December 31, 2022 and Test Set: January 1, 2023, to December 31, 2023. Our experiments integrate financial reports and factor performances of the 50 constituent companies of the SSE 50 Index, providing a comprehensive view of the market. (li2020asset, )The evaluation considers various datasets, including financial reports from the specified periods and performance metrics of different Alpha factors. This extensive dataset allows us to identify causal relationships and patterns between financial features and Alpha factor performances. The diversified portfolio in our evaluation comprises prominent companies within the SSE 50 Index, selected for their extensive market coverage and representation across different sectors of the Chinese A-shares market. This selection underscores the versatility and consistency of our model across various financial environments, showcasing its broad applicability and reliability.
@@ 4.2. Research Questions
Table 3 presents an example combination of 12 alphas generated by our framework, evaluated on the SSE50 constituent stock set. The table details the seed alphas selected by the LLM from each category, along with their respective weights and Information Coefficient (IC) values. The weight combination IC value is quite high as -0.0587. Although some of the seed alphas exhibit relatively low IC values individually, their removal results in a significant drop in the retrained combination weight, indicating their critical role in the overall performance. For example, if we remove alpha #6 the weight combination will drop to -0.055; once we remove the alpha #11, the weight combination will drop to merely 0.0491. This suggests that the seed alpha set selected by the LLM synergizes effectively, providing robust predictive power. (zhang2020information, )
@@ 5.1. Formulaic alphas
Formulaic alphas in quantitative investment refer to systematic, rule-based strategies that generate excess returns by exploiting specific market patterns, anomalies, or inefficiencies. (kakushadze2016101, ) These strategies use predefined mathematical models or algorithms and are often implemented through automated trading systems. The common method for creating formulaic alphas is genetic programming involves structure and numerical mutations to generate new alphas. (cong2021alphaportfolio, ) The gplearn library is enhanced with time-series operators specific forming an alpha-mining framework that incorporates mutual information as the fitness measure. (lin2019alpha, ) AlphaEvolve evolving new alphas from existing ones, using algorithmic graphs for more complex and accurate predictions, though this can reduce interpretability and increase the risk of overfitting. (cui2021alphaevolve, ) Another approach is Machine learning-based alphas are predictive models designed to generate trading signals or investment strategies using advanced machine learning techniques. Key algorithms include neural networks, such as Long Short-Term Memory (LSTM) (hochreiter1997lstm, ) and Transformer models (vaswani2017attention, ), which excel at capturing complex temporal dependencies in time series data. Decision trees, exemplified by models like XGBoost (chen2016xgboost, ) and LightGBM (ke2017lightgbm, ), are valued for their interpretability and ability to handle high-dimensional data. Explainable AI (XAI) (arrieta2020xai, ) methods and tools provide deeper insights into how these models make decisions, fostering greater understanding and trust among investors. Specialized network structures for stock forecasting, such as SFM (zhang2017stock, ), which employs a DFT-like mechanism, have also been developed. More recent research focuses on integrating non-standard data with time series data. For example, REST (xu2021rest, ) combines multi-granular time series data with historical event data to model the overall market. HIST (xu2021hist, ) uses concept graphs alongside regular time series data to capture commonalities in future trends across different stock groups.
@@ 5.3. Multimodal & Multi-Agent LLM
The integration of multimodal large language models (LLMs) in the investment sector has been extensively researched, highlighting their potential to transform financial technology. Li et al. (li2023multimodal, ) demonstrate how multimodal LLMs can process diverse data types such as text, images, and numerical data, offering comprehensive insights and predictive capabilities. Ouyang et al. (ouyang2024modal, ) emphasize the robustness of these models in developing investment strategies that mitigate market volatility risks. By analyzing textual data from news articles and social media, these models can gauge investor emotions and predict market movements with higher accuracy. (zhao2024revolutionizing, ). This is particularly beneficial in high-frequency trading, where anticipating market shifts can provide a competitive edge (zhang2024multimodal, ). In portfolio management, multimodal LLMs offer personalized investment advice tailored to individual risk profiles and financial goals. (li2024cumo, ). These models analyze a client’s financial history and current market conditions to recommend optimal asset allocations and strategies, enhancing client satisfaction and portfolio performance. Additionally, they automate routine tasks such as data entry and report generation, allowing investment professionals to focus on strategic activities (luz2024enhancing, ).
The application of multi-agent LLMs further enhances market analysis and portfolio management. Zhang et al. (zhang2024multimodal, ) highlight how these models leverage vast datasets and advanced natural language processing to interpret financial reports, news, and social media sentiment, providing nuanced insights. Multi-agent systems simulate various market scenarios and stress tests, enabling fund managers to devise robust strategies resilient to market volatility (talebirad2023multiagent, ). The collaborative operation of multiple agents ensures diverse strategies can be tested and optimized in parallel, promoting a balanced approach to risk and return (wang2024rethinking, ). However, the implementation of these advanced models raises ethical and governance considerations. Ensuring transparency, accountability, and adherence to ethical guidelines and regulatory standards is crucial to maintaining investor trust and market integrity. Managing potential biases in data and model training is also essential to prevent systemic risks and ensure fair outcomes (yu2024finmem, ; mundhenk2021symbolic, ).

### v2

@@ 1. Introduction
According to the survey by Grand View Research, in 2023 the global alternative data market was valued at 7.2bn USD and is expected to expand at the compound annual growth rate (CAGR) of 52.1% from 2023 to 2030 representing a complex and dynamic ecosystem that attracts millions of investors seeking profitable investment opportunities(vladi2024quantitative, ). Recent advancements in quantitative trading have significantly enhanced the ability to analyze and process vast amounts of financial data, leading to the development of sophisticated trading strategies. One of the key areas of focus in quantitative trading is Alpha mining—the process of discovering and refining predictive signals, or alpha factors, that forecast asset returns(tulchinsky2015finding, ). Eugene Fama proposed the Efficient Market Hypothesis (EMH) to present that stock price will be the reflection of all the market available information(fama1970efficient, ). However with the market condition change the same information may make different influence on the market, some researchers also found information asymmetry(allen1999genetic, ). Here we found three major challenges to the current Alpha mining process (1) Rigidness of traditional methods: traditional methods for identifying alpha factors in finance often rely on heuristic rules and financial expertise. These rule-based approaches, typically perform well only under specific market conditions and lack the flexibility to adapt to varying market dynamics. (2) Data diversity and integration, recent research has explored the application of machine learning techniques to uncover alpha factors from diverse data sources, such as economic news, social media trends, and historical price movements(cui2021alphaevolve, ; allen1999genetic, ; yu2023generating, ). (3) Adapting to market variability, Market dynamics are fluid, and strategies that perform well in one environment may fail in another. In recent years many researchers proposed to use deep learning methods to predict the market (xu2018stock, ) and form their strategy(chen2019investment, ). Efficiently mining and utilizing alpha factors across different market conditions remains a significant challenge.
@@
To address this challenge, we propose a novel framework that leverages the exploratory power of Large Language Models (LLMs) for Alpha mining and strategy optimization. (1) Flexible Alpha mining with LLMs, the framework starts with the collection and categorization of potential alpha factors from academic research, financial literature, and proprietary datasets. These seed alphas are filtered and categorized using the potential of the LLM, which ensures a comprehensive and diverse set of candidate seed Alphas for further research. All these seed alpha factors are well categorized as momentum factors, fundamental factors, liquidity factors, etc., and based on traditional finance research we believe that these factors are independent of each other(xu2018stock, ). Once new research work is published, our model will be able to incremental update the seed alphas factory. This dynamic approach allows for the discovery of adaptable alpha factors, overcoming the rigidity of traditional methods. (2) Multi-Agent multimodal market evaluation, the seed Alpha factors undergo rigorous back-testing against historical market data to assess their performance. This includes back test under different market conditions to determine which factors can respond to the market for excess returns under different market conditions. To handle the variability in market conditions, we employ a multi-Agent approach. This method allows specialized agents to evaluate the relevance and effectiveness of each alpha factor under multimodal market scenarios. The multi-Agent architecture captures the nuanced dynamics of various market environments from different perspectives. Based on the current market condition the multi-Agent part can filter the out performance seed alphas for us further strategy design. (3) Dynamic strategy optimization, following the evaluation and back-testing phases, a weight gating layer assigns optimal weights to each seed alpha based on current market conditions. This dynamic weighting mechanism ensures that the strategy adapts to changing market environments, optimizing overall performance. The weighted alpha strategies are then aggregated to form a comprehensive investment strategy that maximizes returns while managing risk. Our approach integrates state-of-the-art techniques in machine learning and finance to develop a robust system capable of identifying and optimizing alpha strategies across multiple asset classes. As Figure 7 shows here the work inspired from the idea in real-life quant company practice, for the real quant investment company they generally have four major departments. The first one is the research department’s response to find the valuable trading signals and record them; the second part is the trading department need to analyze the current market condition and select the ideal alphas for future portfolio construction; the third part is the risk management part, they need to evaluate and make the strategy suitable for the risk preference; the last part will be the portfolio manager, they mix up the traders idea construct the final trading strategy for profit(fabozzi2010quantitative, ).
@@ 2.2. Seed Alpha Mining & Selection
In quantitative trading, the static nature of traditional heuristic and rule-based alpha mining methods poses significant limitations (e.g. Information coefficient), particularly in the context of rapidly changing market conditions(zhang2020information, ). These methods often fail to adapt dynamically, leading to suboptimal performance and the selected Alpha may loss function. Moreover, the reliance on structured financial data, such as historical prices and fundamental indicators, overlooks the rich insights that can be derived from alternative data sources like economic news and social media trends. To address these challenges, we propose a novel framework that leverages the exploratory power of Large Language Models (LLMs) and a multi-agent system for seed alpha mining and selection. Our approach integrates state-of-the-art machine learning techniques with financial principles to create a robust system capable of dynamically adjusting to market variability and optimizing alpha strategies across $k$ categories of seed alphas. This multimodal framework links current market conditions with alpha back-test results, allowing LLMs to identify the most suitable seed alphas while assigning confidence scores $\theta$ to ensure reliability. Specialized agents further customize the evaluation based on risk preferences and expertise, ensuring a comprehensive and adaptive solution for maximizing returns and managing risks in diverse market environments.
@@ 3.3. Multimodal & multi-Agent Evaluate Alpha set
The multi-agent system comprises various agents, each designed with distinct risk preferences and investment strategies(luo2025large, ). These agents analyze the multimodal data and evaluate the seed alphas based on their respective criteria. Each agent conducts a risk preference analysis, assessing the seed alphas according to its predefined risk parameters. This ensures that the evaluation covers a wide spectrum of risk perspectives. Agents then select alpha factors that align with their risk preferences and market analyses, ensuring relevance and robustness under different market scenarios. To enhance the reliability of the selected alpha factors, a confidence scoring mechanism is implemented. Each agent assigns a confidence score to the alpha factors based on their analysis and historical performance, reflecting the agent’s assessment of the alpha factor’s predictive power and robustness. Selected alpha factors undergo rigorous backtesting using historical market data to evaluate their performance under various market conditions. This step is crucial for identifying the most promising alpha factors for the final strategy. Key evaluation metrics include performance indicators such as the Information Coefficient (IC) and Sharpe Ratio etc(sharpe1994sharpe, ). Additionally, seed alpha factors from each category must meet specific confidence score thresholds to be considered for the final selection, ensuring that only the most reliable and robust alpha factors are chosen.
@@ 3.4. Optimal weight alpha strategy
To optimize the weights of selected seed alphas, we employ a Deep Neural Network (DNN) to fit alpha values to future yields. The network architecture comprises three layers: an input layer, a hidden layer, and an output layer. The input layer takes daily alpha calculations derived from historical closing prices as input features. The hidden layer, equipped with ten nodes, utilizes ReLU activation functions to introduce non-linearity and enhance the model’s learning capability(chen2020dynamic, ). The output layer, consisting of a single node, predicts future yields. During the training phase, the network employs backpropagation and gradient descent to minimize the loss function, which quantifies the discrepancy between predicted and actual yields. This iterative process adjusts the network’s weights and biases to improve prediction accuracy. To ensure the model’s generalizability and prevent overfitting, a separate validation set is utilized. This methodology establishes a robust framework for predicting future yields based on historical alpha values, thus forming an effective investment strategy. The DNN model processes the input data through the hidden layer, transforming it with learned weights and biases. The final output is generated by applying another set of weights, biases, and an activation function in the output layer. Our approach demonstrates the efficacy of leveraging deep learning techniques to optimize alpha weights, thereby enhancing the predictive performance of investment strategies.
@@ 4.1. Datasets
Our study focuses on financial data from the Chinese A-shares market, specifically targeting the SSE 50 Index. The dataset encompasses six primary features as original inputs for our Alpha factors: open, close, high, low, volume, and volume-weighted average price (VWAP). To ensure rigorous evaluation and robust model performance, the dataset is divided into distinct time periods: Training Set: January 1, 2021, to December 31, 2022 and Test Set: January 1, 2023, to December 31, 2023. Our experiments integrate financial reports and factor performances of the 50 constituent companies of the SSE 50 Index, providing a comprehensive view of the market(li2020asset, ). The evaluation considers various datasets, including financial reports from the specified periods and performance metrics of different Alpha factors. This extensive dataset allows us to identify causal relationships and patterns between financial features and Alpha factor performances. The diversified portfolio in our evaluation comprises prominent companies within the SSE 50 Index, selected for their extensive market coverage and representation across different sectors of the Chinese A-shares market. This selection underscores the versatility and consistency of our model across various financial environments, showcasing its broad applicability and reliability.
@@ 4.2. Research Questions
Table 3 presents an example combination of 12 alphas generated by our framework, evaluated on the SSE50 constituent stock set. The table details the seed alphas selected by the LLM from each category, along with their respective weights and Information Coefficient (IC) values. The weight combination IC value is quite high as -0.0587. Although some of the seed alphas exhibit relatively low IC values individually, their removal results in a significant drop in the retrained combination weight, indicating their critical role in the overall performance. For example, if we remove alpha #6 the weight combination will drop to -0.055; once we remove the alpha #11, the weight combination will drop to merely 0.0491. This suggests that the seed alpha set selected by the LLM synergizes effectively, providing robust predictive power(zhang2020information, ).
@@ 5.1. Formulaic alphas
Formulaic alphas in quantitative investment refer to systematic, rule-based strategies that generate excess returns by exploiting specific market patterns, anomalies, or inefficiencies(kakushadze2016101, ). These strategies use predefined mathematical models or algorithms and are often implemented through automated trading systems. The common method for creating formulaic alphas is genetic programming involves structure and numerical mutations to generate new alphas(cong2021alphaportfolio, ). The gplearn library is enhanced with time-series operators specific forming an alpha-mining framework that incorporates mutual information as the fitness measure(lin2019alpha, ). AlphaEvolve evolving new alphas from existing ones, using algorithmic graphs for more complex and accurate predictions, though this can reduce interpretability and increase the risk of overfitting(cui2021alphaevolve, ). Another approach is Machine learning-based alphas are predictive models designed to generate trading signals or investment strategies using advanced machine learning techniques. Key algorithms include neural networks, such as Long Short-Term Memory (LSTM) (hochreiter1997lstm, ) and Transformer models (vaswani2017attention, ), which excel at capturing complex temporal dependencies in time series data. Decision trees, exemplified by models like XGBoost (chen2016xgboost, ) and LightGBM (ke2017lightgbm, ), are valued for their interpretability and ability to handle high-dimensional data. Explainable AI (XAI) (arrieta2020xai, ) methods and tools provide deeper insights into how these models make decisions, fostering greater understanding and trust among investors. Specialized network structures for stock forecasting, such as SFM (zhang2017stock, ), which employs a DFT-like mechanism, have also been developed. More recent research focuses on integrating non-standard data with time series data. For example, REST (xu2021rest, ) combines multi-granular time series data with historical event data to model the overall market. HIST (xu2021hist, ) uses concept graphs alongside regular time series data to capture commonalities in future trends across different stock groups.
@@ 5.3. Multimodal & Multi-Agent LLM
The integration of multimodal large language models (LLMs) in the investment sector has been extensively researched, highlighting their potential to transform financial technology. Li et al(li2023multimodal, ). demonstrate how multimodal LLMs can process diverse data types such as text, images, and numerical data, offering comprehensive insights and predictive capabilities. Ouyang et al(ouyang2024modal, ). emphasize the robustness of these models in developing investment strategies that mitigate market volatility risks. By analyzing textual data from news articles and social media, these models can gauge investor emotions and predict market movements with higher accuracy (zhao2024revolutionizing, ). This is particularly beneficial in high-frequency trading, where anticipating market shifts can provide a competitive edge (zhang2024multimodal, ). In portfolio management, multimodal LLMs offer personalized investment advice tailored to individual risk profiles and financial goals (li2024cumo, ). These models analyze a client’s financial history and current market conditions to recommend optimal asset allocations and strategies, enhancing client satisfaction and portfolio performance. Additionally, they automate routine tasks such as data entry and report generation, allowing investment professionals to focus on strategic activities (luz2024enhancing, ).
The application of multi-agent LLMs further enhances market analysis and portfolio management. Zhang et al (zhang2024multimodal, ). highlight how these models leverage vast datasets and advanced natural language processing to interpret financial reports, news, and social media sentiment, providing nuanced insights. Multi-agent systems simulate various market scenarios and stress tests, enabling fund managers to devise robust strategies resilient to market volatility (talebirad2023multiagent, ). The collaborative operation of multiple agents ensures diverse strategies can be tested and optimized in parallel, promoting a balanced approach to risk and return (wang2024rethinking, ). However, the implementation of these advanced models raises ethical and governance considerations. Ensuring transparency, accountability, and adherence to ethical guidelines and regulatory standards is crucial to maintaining investor trust and market integrity. Managing potential biases in data and model training is also essential to prevent systemic risks and ensure fair outcomes (yu2024finmem, ; mundhenk2021symbolic, ).